OpenStudy (anonymous):

g(x) = 3(x − 2), find the value of x if g(x) = 6 Explain please =)

jimthompson5910 (jim_thompson5910):

g(x) = 3(x − 2) 6 = 3(x − 2) ... Plug in g(x) = 6 Now solve for x

OpenStudy (anonymous):

that doesnt make sence @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I'm just replacing g(x) with 6

jimthompson5910 (jim_thompson5910):

That's all I did. You need to solve for x to find the answer

OpenStudy (anonymous):

oh! ok thank you